Required information [The following information applies to the questions displayed below.] The following information relates to a joint production process for three products, with a total joint production cost of $100,000. There are no separable processing costs for any of the three products. Product Sales Value at Split-Off Units at Split-Off 1 $ 130,000 240 2 50,000 960 3 20,000 1,200 $ 200,000 2,400 What percentage of joint cost is allocated to each of the three products using the physical units method?
Answer #1

Calculate percentage of joint cost allocated to three products using physical units method

Product Units Joint Cost allocated Percentage
1 240 240/2400 = 10%
2 960 960/2400 = 40%
3 1200 1200/2400 = 50%
Total 2400 100%